OpenStudy (anonymous):

Whats the greatest common factor of 8 and 8?

OpenStudy (compassionate):

What are the two greatest numbers that can be multiplied to get 8?

OpenStudy (anonymous):

uh.. 4?

OpenStudy (anonymous):

\[1 \times 8 = 8\] \[2 \times 4 = 8\] \[4 \times 2 = 8\] \[8 \times 1 = 8\]

OpenStudy (anonymous):

I'm a bit confused... I thought it was suppose to be ONE number.

OpenStudy (anonymous):

factors of 8: 1, 2, 4, 8

OpenStudy (anonymous):

So Whats the greatest common factor?

OpenStudy (anonymous):

8 i guess

OpenStudy (anonymous):

right?